AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 189     Entry time: Tue Feb 8 08:13:53 2022
Author: OH, CB, DF 
Subject: Hopefully effect of beam tuning on detectors 
09:00 Detector still biased to 60V and leakage current has been mostly stable overnight.
      Only minor fluctuations observed - attachment 1

      FEE temps all ok - attachment 2
      Stats at 60V - attachment 3
      System wide checks all ok

      Increasing bias to 80V
      60V 9.45uA
      70V 9.7uA
      80V 80 9.82uA Still have diode response

10:15 In prep for taking data with beam tuning have tested the merger and tapeserver.
      To enable writing to file had to implement the method in https://elog.ph.ed.ac.uk/DESPEC/36
      Probably a corrupution caused in the difference between the carme system and the aida system.
      Now can write data to file. To limit the data of FEEs 01-04 have set threshold to 0xFF

10:33 Ion source was started this morning but there was no noticable effect on the leakage currents.

10:56 Leakage current is slowly rising continuing the trend at 60V. - Attachment 4
      This could just be the temperature shift during the day or linked to the ion source being powered on.

11:18 Pulser peaks for the FEEs using layout 3 and 4.
      Layout 3 - attachment 5
      Layout 4 - attachment 6
      The pulser peaks appear to be asymetric and in FEE 7 we are observing a double peak.
      Will check the circuit when we next have access.

11:23 Have started writing data to disk in advance of them putting beam into the ring.
      Writing to R4_8 at around 1.8MB/s. N.B. TapeServer shows the same behavior as AIDA with writing nothing to files R4_0->R4_7
      FEE 01-04 slow comp 0xFF
      FEE 05-08 slow comp 0x64
      Pulser running at 50Hz or so

14:50 Having been writing to DAQ since lunch.
      Leakage currents were mostly stable but we are now seeing an increase in leakage currents at a much greater rate - attachment 7

14:58 Attachment8 Updated leakage currents.

15:00 Due to the rapidly increasing bias we asked them to stop the beam.
      Motors were then cycled but we did not notice the leakage coming down.
      Bias was turned off and then motors cycled again.
      Bias was then set to 60V - Run R5
      Bias raised to 80V - R6
      Carlo requested we stop writing data to disk

16:25 Stopped writing to disk. Will retry bias at 80V and monitor.

18:13 Motors cycled before leaving for night no beam in ring - attachment 9
      to test if the increase is beam related we will leave at 80V overnight with the beam blocked all night.
      To avoid the rapidly increasing leakage current the iSet on the CAEN module is at 11uA with a trip time of 2s.
      
Attachment 1: 2022-02-08-09-15-48-BiasOvernight.png  8 kB  | Hide | Hide all
2022-02-08-09-15-48-BiasOvernight.png
Attachment 2: 2022-02-08-09-25-23_FEE_TEMP.png  97 kB  | Hide | Hide all
2022-02-08-09-25-23_FEE_TEMP.png
Attachment 3: 220208_0926_Stats.png  75 kB  | Hide | Hide all
220208_0926_Stats.png
Attachment 4: 220208_1056_Bias.png  8 kB  Uploaded Tue Feb 8 09:57:38 2022  | Hide | Hide all
220208_1056_Bias.png
Attachment 5: 220208_1118_Layout3.png  176 kB  Uploaded Tue Feb 8 10:20:35 2022  | Hide | Hide all
220208_1118_Layout3.png
Attachment 6: 220208_1118_Layout4.png  185 kB  Uploaded Tue Feb 8 10:20:35 2022  | Hide | Hide all
220208_1118_Layout4.png
Attachment 7: 220208_1435_Bias.png  19 kB  Uploaded Tue Feb 8 13:51:14 2022  | Hide | Hide all
220208_1435_Bias.png
Attachment 8: 220208_1458_Bias.png  18 kB  Uploaded Tue Feb 8 13:58:59 2022  | Hide | Hide all
220208_1458_Bias.png
Attachment 9: 220208_1824_Bias_shake.png  19 kB  Uploaded Tue Feb 8 17:25:16 2022  | Hide | Hide all
220208_1824_Bias_shake.png
ELOG V3.1.4-unknown